# Towards Practical Finite Sample Bounds for Motion Planning in TAMP

### v1

@@ 4 An Attempt to Extend to KNN PRMs
Let $X_{free}$ be a configuration space, and let $N\subset X_{free}$ be a set of samples. Suppose that $G=(V,E)$ is a $K$-nearest neighbor roadmap graph on $N$.
We say $r\in\operatorname{\mathbb{R}}^{+}$ is a connection radius of $G$ if $\forall u,v\in V$ if $\left\lVert u-v\right\rVert\leq r$, then $u$ is at most the $K$th nearest neighbor of $v$, or $v$ is at most the $K$th nearest neighbor of $u$.
@@
We also note that Eq. 8 is a conservative approximation for the true connectivity of the KNN PRM (Fig. 3a).
The upper bound in Eq. 8, as a function of a random geometric graph of degree $k$, is a random variable, so we must understand its behavior before we use it as a connection radius as in the analysis for radius PRMs. Fortunately, as the number of samples increases, the realizations of the upper bound tend to concentrate around similar values222 The right-hand quantity in Eq. 8 obeys concentration of measure phenomena, e.g. a generalization of how sums of i.i.d. random variables tend to only deviate from their means by small amounts as stated in the Central Limit Theorem [35]. . We will exploit this fact to provide a conservative lower bound to the expression in Eq. 8 to study how the connection radii of a KNN PRM shrinks as the number of samples increases.
@@
Let $V$ be a set of $n$ i.i.d. uniform samples from collision-free configuration space $X_{free}$ used to build a $K$-nearest neighbor PRM, and suppose that $18K<n-$ 1.333The reason we state this mild technical condition (it is common to sample far more than the number of neighbors) is explained in the appendix. Then
@@
with probability at least $1-\gamma$.
We now use the lower bound stated in Eq. 9 to stand in for the connection radius assumption in the statement of Lemma 1:
@@
Let $V$ be a set of $n$ i.i.d. uniform samples from the collision-free configuration space $X_{free}$ used to build a $K$-nearest neighbor PRM, and suppose that $18K<n-$ 1. Let $r\in\operatorname{\mathbb{R}}^{+}$. If:
$r\leq\left[\frac{K-\sqrt{2K\log(n/\gamma)}}{(n-1)\operatorname{\mathrm{P}}(B_{1}^{d})}\right]^{1/d},$
then $r$ is a connection radius with probability at least $1-\gamma$.
@@
  • As the failure probability $\gamma$ increases, the effective connection radius grows on the order of $O\left(\frac{K\log(1/\gamma)}{\operatorname{\mathrm{P}}(B_{1}^{d})}\right)^{1/2d}$.
  • As the number of samples $n$ increases, the effective connection radius decays on the order of $o\left(\frac{K}{n\cdot\operatorname{\mathrm{P}}(B_{1}^{d})}\right)^{1/d}$.
We suspect that the conservative $\sqrt{\log n}$ factor (which makes the bound vacuous as $n\to\infty$ and the culprit for the small $o$ above) can be improved with a tighter analysis. However, for small $n$ (less than 1,000,000), we empirically find that the bound is fairly tight (numerical experimental results are included in the appendix). Furthermore, we can even define a similar numerical algorithm to Alg. 2 that can search for a tighter radius as a function of the number neighbors, samples, and failure probability in the appendix.
@@
Let $X_{free}$ be the collision-free configuration space, and suppose Near is defined be a $K$-nearest neighbor search.
Suppose $N\subset\mathcal{X}$ be a set of $n$ i.i.d. uniform random samples. If:
$n\geq\max\left\{\frac{4}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{4}{\gamma},\frac{8d}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log\frac{13}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\right\},$ (10)
@@
then for all $x_{s},x_{g}\in X$, if there exists a solution path between $x_{s},x_{g}$ that has clearance $2\alpha$, then the PRM will return a feasible path from $x_{s}$ to $x_{g}$ with probability $1-\gamma$.
Figure 3: Left: A depiction of the effective connection radius for a PRM graph with $K=1$. Note that we have a conservative approximation that fails to capture ‘long distance edges’ between isolated vertices. Right: A plot for the ratio of the connection radius to the net radius as a function of the number of sampled vertices with a fixed failure probability of $\gamma=0.1$ and $K=256$ (see Theorem 3.1). The conditions for Theorem 3.1 are satisfied for all samples which have ratios above 4.0 (dotted green line). All of the curves are below the line, so the conditions of Theorem 4.1 are not satisfied for any realistic $n$.
Equation 10 expresses the same sample condition from Lemma 2, and Equation 11 fulfills the connection radius condition stated in the hypothesis of Lemma 1. Unfortunately, as shown by Fig. 3, these two conditions can never be satisfied at the same time. The effective connection radius decays at a rate faster than $o(1/n^{1/d})$, while the largest net radius to be found after $n$ samples is $\Omega(1/n^{1/d})$. Lemma 1 stipulates that these two quantities must form a net-radius to connection-radius ratio of 1:4, which is never true (Fig. 3).
We suspect the vacuousness of the bound comes from two issues: looseness in the analysis of the effective radius and looseness in the union bound combining Prop. 1 and Theorem 3.1. Improvement in both steps of the argument may result in a sample bound for KNN PRMs.
@@ 5 Experiments
We seek to experimentally verify the tightness of our sampling bound of convergence to path feasibility by Algorithm 2. Since the conditions on Theorem 4.1 prevent us from formally bounding the number of samples required by a KNN PRM, we explore using the radius bound (Theorem 3.1 and Alg. 2) as a possible sampling heuristic for KNN PRMs, since KNN connection approach is better suited to higher-dimensional problems with many samples.
@@ 5.1 Bound Tightness in a Narrow Passageway Problem
Our clearance experiments are conducted in a ‘narrow hallway’ problem that can be generalized to arbitrarily high dimensions. The configuration space can be characterized by
@@
The hallway is the narrow passage and its center path $[-0.5,0.5]\times\{0\}^{d-1}$ has clearance $\delta.$ We construct many instances of this hallway problem for varying $\delta<0.5$ and for $d=\{2,3,6,7,14,20\}$. We form a Monte Carlo estimate of the probability of finding a path in the narrow passages by building 100 PRMs and checking if they return a solution to the query $x_{s}=(-0.5,0,\dots,0),x_{g}=(0.5,0,\dots,0)$ for samples $n=100,1000,10000$. We then compare the number of samples the numerical bound (Alg. 2) predicts to find a path in the passage with the Monte Carlo estimated probability. To explore the use of Alg. 2 as a sampling heuristic for KNN PRMs, we compute the bound according to the clearance of the environment and set the probability of failure to $\gamma=0.99$. The results are in Table 1.
@@
[…]
For radius PRMs, we find that the bound is loose 2-3 orders of magnitude for the planar hallway problem and becomes looser as the number of dimensions increases. In KNN PRMs, the radius bound (as a sampling heuristic) is tight for 1-2 orders of magnitude and loosens as the number of dimensions grows. This means other conditions besides the sufficient conditions outlined in Lemma 1 can allow the PRM to find a path through the narrow passage. Surprisingly, we find that Alg. 2 is better suited as a sampling heuristic in lower dimensions for KNN PRMs than for radius PRMs.
@@ 5.2 TAMP Experiments
Our experiments in Table 1 suggest that the radius PRM bound may be a useful sampling heuristic for a KNN PRM, which is more tractable to compute in practice. We now aim to assess the usefulness of such a bound in the context of a TAMP planner that must solve multiple motion planning subproblems.
@@
(a) Our 2DoF mobile-base manipulation problem setting (left) and a scatter plot showing TAMP success vs planning time (right) for the three PRM-based motion planning methods. Each point on the scatter plot is the performance and planning time across 100 environment randomizations in which we vary the number and size of objects. The iPRM and sPRM points represent different settings of N.
(b) Our 6DoF peg-in-hole problem setting (left) and corresponding baseline evaluations (right).
@@
We examine the bound’s performance in a planar task involving a circular robot with 2DoF and its scalability to higher dimensions as we experiment with a 6DoF peg-in-hole insertion task. In the task shown in Figure 4(a), a mobile-base robot must transfer all yellow objects (5-10) of varying sizes from one room to another by passing through a narrow passage. All objects are made to be small enough to fit through the hallway. In Figure 4(b), a 6DoF KUKA robot arm must move all objects (5-10) of varying size from an initial region into a packing container on the right. Each robot’s planner must solve a sequence of motion planning subproblems of varying difficulty based on the size of the object the robot is holding. We make use of the Focused PDDLStream TAMP algorithm [9], which proposes high-level mode sequences (e.g., object grasps and releases) and calls a motion planning subprocess to refine the task skeleton or deem it infeasible.
@@
Using an adaptive bound selection strategy, which we will call aPRM, we compute an upper bound for the number of samples PRM uses upon request from the TAMP solver. Given the size of the object being held, we are able to compute $\delta$ by subtracting the hallway size from the extent of the robot when holding the object, which is sometimes larger than the extent of the robot itself. We then call sPRM outlined by Karaman and Frazzoli [15] (see appendix) with an $N$ determined by the numerical bound for radius PRMs (Alg. 2).
@@
We compare this strategy to two other baseline strategies. First, we compare to the sPRM motion planner with a fixed N selected by hand and fixed across subproblems, which is standard practice in TAMP. Second, we compare to an incremental version of sPRM, which we call iPRM. iPRM iteratively calls sPRM with an increasing max sample bound $K$, which increases geometrically according to $M_{i+1}=cM_{i}$ until a motion plan is found or $M\geq{N}$. Our experiments use $c=1.1$. We compare both of these strategies with aPRM in Figure 4(a).
@@
We find that in the planar problem, we observe that aPRM is able to solve nearly all problems with less computation time than sPRM and iPRM. The radius numerical bound helps the planner effectively trade sampling effort with the complexity of the motion planning problems, whereas sPRM either undersamples or oversamples problems with a single threshold, and iPRM wastes computation time building up to the correct number of samples. In the 6DoF problem, our results are weaker, where some instances of iPRM scale up to a sufficient number of samples faster than aPRM takes in a single computation. This is consistent with our observation that the numerical bound becomes looser as the number of dimensions increases.
@@ 6 Related work
Another approach to terminating sampling-based motion planners is to decide the motion planning problem’s feasibility directly. In the literature, there are two common classes of methods to solve this problem. One common approach is to show the existence of a ‘disconnection proof,’ an obstacle region that fully encloses the goal from an outer starting configuration [1, 18, 19]. These proofs usually consist of discrete representations of hypersurfaces that fully enclose the region around the goal. Others used volumetric approximations of the configuration space or obstacle region [33, 23, 36] to find a path or refute feasibility (up to the accuracy of the approximation). These methods scale poorly with a dimension of the configuration space and often empirically are intractable in larger than four dimensions. These representations are computationally expensive to construct, and in most situations, evaluating a cheaper probabilistic bound may be more desirable.
@@
Much of our understanding of PRMs with random sampling is from asymptotic results stated in the limit of infinite samples. Karaman and Frazzoli [15] proved that paths returned by PRM (and the more space-frugal PRM*) are both asymptotically optimal. Tighter estimates on the necessary connection radius to ensure graph connectivity have been obtained as well [28]. All of these results are proven using tail-field type arguments that are difficult to extend to a finite number of samples.
@@
In addition to Hsu et al.’s work on configuration space expansiveness [11], there are several instances of finite-random sample bounds. Kavraki et al. [16] and Dobson et al. [7] gave a bound on the number of samples required to return a path and converge to a near-optimal path, respectively, by assuming and clearance and path length. They both make stronger assumptions on the solution, like path length, for simpler arguments that do not rely on covering sets. These assumptions also preclude them from application in scenarios where the knowledge of the solution path length is unknown. Deterministic low-discrepancy and low-dispersion sampling processes are sample sets that cover the space with a prespecified density. They are justified by superior bounds on sample dispersion (largest unsampled ball) as number of samples increases [25] and present better empirical performance in PRMs than random samples [14]. However, their performance gains over random samples degrades as dimensions increase [13], and while incremental sequences exist, the best-performing sequences are non-incremental [25, 21]. There are techniques that layer varying-density deterministic sample sets to produce an incremental sequence [20], but at the cost of repeating computational effort from coarser grids. For these reasons, random samples are a suitable choice for our problem setting.
@@
Outside of motion planning theory, there are sample bounds for random coverings have been studied in discrete geometry and computational learning theory [22]. The techniques to prove bounds of these kinds has been credited to Vapnik and Chernonenkis [32], Sauer [26], and Shelah [27], who sought to find small coverings by the probabilistic method [^4] Please see the appendix for more details of how these details fit into the proofs of our results.. In computational learning theory, Blumer et al. estimate the probability that the samples form a covering [2] to form probably approximate correct learning guarantees. Their results for the smallest $\alpha$-net (right-hand argument in Lemma 2) were confirmed to be tight up to constants by Kómlos et al. [17] by a matching lower bound. This paper draws a connection between these results and motion planning.
@@
Lastly, this paper shows the potential usefulness of a sampling bound in a search-then-sample style planner. Search-then-sample is a fairly common approach to TAMP [9, 5, 6]. Such planners use sampling-based motion planners use ad hoc sample thresholds or runtimes, which we show can be replaced with a sample bound. In addition to setting resource limits on individual motion planning subproblems, a sample bound could be used to guide the task-level search as a heuristic to decide which motion planning problems to attempt. Some related works along these lines include Cambon et al., wherein the number of failed queries is a heuristic to decide which task plans to prioritize for motion computation [4] and Srivastava et al., where estimated subproblem feasibility under different object pose configurations is used to discover obstacles that must be moved to construct a multi-step plan [29]. We do note that some methods do not directly benefit from a sample bound by the nature of their construction. Some TAMP planners explore by constructing motions first [8, 34], and only make task decisions based on constructed motion plans. Others forgo sampling-based motion planning completely in favor of optimization-based obstacle avoidance [30].

### v3

@@ 4 An Attempt to Extend to KNN PRMs
Let $X_{free}$ be a configuration space, and let $N\subset X_{free}$ be a set of samples. Suppose that $G=(V,E)$ is a $K$-nearest neighbor roadmap graph on $N$. We say $r\in\operatorname{\mathbb{R}}^{+}$ is a connection radius of $G$ if $\forall u,v\in V$ if $\left\lVert u-v\right\rVert\leq r$, then $u$ is at most the $K$th nearest neighbor of $v$, or $v$ is at most the $K$th nearest neighbor of $u$.
@@
We observe that Eq. 8 is a conservative approximation for the true connectivity of the KNN PRM (Fig. 3a), and thus, so is our conception of the effective connection radius.
The upper bound in Eq. 8, as a function of a random geometric graph of degree $k$, is a random variable, so we must understand its behavior before we treat it as a connection radius. Fortunately, as the number of samples increases, the realizations of the upper bound tend to concentrate around similar values. We exploit this fact to prove a lower bound to the expression in Eq. 8 to study how the connection radii of a KNN PRM shrinks as the sample count increases.
@@
Let $V$ be a set of $n$ i.i.d. uniform samples from collision-free configuration space $X_{free}$ used to build a $K$-nearest neighbor PRM. Then with probability at least $1-\gamma$:
@@
We now use the lower bound stated in Eq. 9 to derive a conservative approximation of the effective connection radius using Lemma 3:
@@
Let $V$ be a set of $n$ i.i.d. uniform samples from the collision-free configuration space $X_{free}$ used to build a $K$-nearest neighbor PRM. Let $r\in\operatorname{\mathbb{R}}^{+}$. Then $r$ is a connection radius with probability at least $1-\gamma$ if
$r\leq\left[\frac{K-\sqrt{2K\log(n/\gamma)}}{(n-1)\operatorname{\mathrm{P}}(B_{1}^{d})}\right]^{1/d}.$
@@
  • As $\gamma$ increases, the effective connection radius grows by $O\left(\frac{K\log(1/\gamma)}{\operatorname{\mathrm{P}}(B_{1}^{d})}\right)^{1/2d}$.
  • As $n$ increases, the effective connection radius decays by $o\left(\frac{K}{n\cdot\operatorname{\mathrm{P}}(B_{1}^{d})}\right)^{1/d}$.
We suspect that the conservative $\sqrt{\log n}$ factor (which makes the bound vacuous as $n\to\infty$ and the reason for the small $o$ above) can be improved with a tighter analysis. We can even define a similar numerical algorithm to Alg. 2 that can search for a tighter radius as a function of the number of neighbors, samples, and failure probability. For small $n$ (less than 1,000,000), we empirically find that a radius estimated by the numerical algorithm is reasonably tight (numerical experimental results are included in the supplemental material).
@@
Suppose $N\subset X_{free}$ be a set of $n$ i.i.d. uniform random samples, from which we build a $K$-nearest neighbor roadmap. If:
$n\geq\max\left\{\frac{4}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log_{2}\frac{4}{\gamma},\frac{8d}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\log_{2}\frac{13}{\operatorname{\mathrm{P}}(B^{d}_{\alpha})}\right\},$ (10)
@@
then for all $x_{s},x_{g}\in X_{free}$, if there exists a solution path between $x_{s},x_{g}$ that has clearance $2\alpha$, then the PRM will return a feasible path from $x_{s}$ to $x_{g}$ with probability $1-\gamma$.
Figure 3: Left: A depiction of the effective connection radius for a PRM graph with $K=1$. Note that we have a conservative approximation that fails to capture ‘long distance edges’ between isolated vertices. Right: A plot of the ratio of the connection radius to the net radius as a function of the number of sampled vertices with a fixed failure probability of $\gamma=0.1$ and $K=256$, as determined by their numerical procedures. The conditions for Theorem 3.1 are satisfied for all sample counts with ratios above 4.0 (dotted red line). All curves are below the line, so the conditions of Theorem 4.1 are not satisfied for any realistic $n$.
Equation 10 expresses the same sample condition from Lemma 2, and Equation 11 fulfills the connection radius condition stated in Lemma 1. Unfortunately, as shown by Fig. 3, these two conditions can never be satisfied at the same time. The effective connection radius decays at a rate faster than $o(1/n^{1/d})$, while the largest net radius to be found after $n$ samples is $\Omega(1/n^{1/d})$. Lemma 1 stipulates that these two quantities must form a net-radius to the connection-radius ratio of 1:4, which is not true. A less conservative estimate of the ratio using the corresponding numerical algorithms for each radius does not meet this requirement either (Fig. 3). We suspect the vacuousness of the bound comes from two issues: looseness in the analysis of the effective radius and looseness in the union bound combining Prop. 1 and Theorem 3.1. Improvement in both steps of the argument may result in a non-vacuous sample bound for KNN PRMs.
@@ 5 Experiments
We seek to experimentally verify the tightness of Algorithm 2 to estimate the number of samples sufficient to find a path. In practice, the KNN connection approach is better suited to higher-dimensional problems with many samples. Since the conditions on Theorem 4.1 prevent us from formally bounding the number of samples required by a KNN PRM, we explore using the radius numerical algorithm as a possible sampling heuristic to guide the number of samples chosen to build KNN PRMs.
@@ 5.1 Bound Tightness in a Narrow Passageway Problem
Our clearance experiments are conducted in a ‘narrow hallway’ problem that can be generalized to arbitrary dimensions. We take the configuration space to be a union of boxes:
@@
The hallway is the narrow passage and its center path $[-0.5,0.5]\times\{0\}^{d-1}$ has clearance $\delta.$ We construct many instances of this hallway problem for varying $\delta<0.5$ and for $d=\{2,3,6,7,14,20\}$. We form a Monte Carlo estimate of the probability of finding a path in the narrow passage by running 100 PRMs and checking the query $x_{s}=(-0.5,0,\dots,0),x_{g}=(0.5,0,\dots,0)$ for samples $n=100,1000$. We then compare the number of samples predicted by the numerical bound (Alg. 2) with the estimated probability. To explore the use of Alg. 2 as a sampling heuristic for KNN PRMs, we compute the bound according to the clearance of the environment and set the probability of failure to $\gamma=0.01$.
Table 1: A comparison of the empirical performance of the radius and KNN PRM ($K=32$) and the numerical bound for the hallway problem with $\delta=0.499,0.25,0.125$, $d=2,3,4,5,6$, and samples sets with $n=100,1000$. The left side of the slash is an MC estimate of the prob. of finding a solution. The right side is the number of samples as output by Alg. 2 using the estimated prob. (minus a machine epsilon to return finite samples) for radius PRMs or $\gamma=0.01$ for the KNN PRM heuristic. If all MC iterations for $n=100$ returned a path, the corresponding box for $n=1000$ is redundant and replaced with a hyphen.
@@
Our results can be found in Table 1. For radius PRMs, the bound predicts two to three orders of magnitude more than needed for the planar hallway problem and becomes progressively looser as the number of dimensions increases. This means other conditions besides the sufficient conditions outlined in Lemma 1 can allow the PRM to find a path through the narrow passage. In KNN PRMs, the radius bound (as a sampling heuristic) is useful for 1-2 orders of magnitude for low dimensions ($d=2,3$) and loses utility as the number of dimensions grows. Interestingly, we find that Alg. 2 is better suited as a sampling heuristic in lower dimensions for KNN PRMs than as a guarantee for radius PRMs.
We do not find the looseness of the bound surprising. As the number of dimensions increases, so does the volume of the hallway ‘ends.’ The bound applies when the entire volume of the configuration space is covered. Since the hallway ends do not need to be densely sampled to yield a path, the bound becomes increasingly loose as the number of dimensions increases.
@@ 5.2 TAMP Experiments
Our experiments in Table 1 suggest that the radius PRM bound may be a useful sampling heuristic for a KNN PRM, which is more tractable to compute in practice. We now aim to assess the usefulness of the numerical algorithm in the context of a TAMP planner that must solve multiple motion planning subproblems in scenarios where the clearance of the narrow passage is known (Sec. 2 discusses the more general case when the path clearance is unknown).
@@
(a) Our 2DoF mobile-base manipulation problem setting (left) and a scatter plot showing TAMP success vs planning time (right) for the three PRM-based motion planning methods. Each point on the scatter plot is the performance and planning time across 100 environment randomizations in which we vary the number and size of objects. The iPRM and sPRM points represent different maximum number of samples $N$.
(b) Our 6DoF manipulation problem setting (left) and corresponding baselines (right).
@@
We examine the bound’s performance in a planar task involving a circular robot with 2DoF and a 6DoF peg-in-hole insertion task. In the task shown in Figure 4(a), the robot must transfer all yellow objects (5-10) of varying sizes from one room to another by passing through a narrow passage. All objects are made to be small enough to fit through the hallway. In Figure 4(b), a 6DoF KUKA robot arm must move all objects (5-10) of varying size from an initial region into a packing container. Each robot’s planner must solve a sequence of motion planning subproblems of varying difficulty based on the size of the object the robot is holding. We use the Focused PDDLStream TAMP algorithm [9], which proposes plan skeletons consisting of high-level mode sequences (e.g., object grasps and releases) and calls a PRM to refine each candidate task skeleton.
@@
Using an adaptive bound selection strategy, which we will call aPRM, we compute an upper bound for the number of samples the KNN PRM uses upon request from the TAMP solver. Given the size of the held object, we can compute $\delta$ by subtracting the hallway size from the extent of the robot when holding the object, which can be larger than the extent of the robot itself. We then call the KNN version of sPRM outlined by Karaman and Frazzoli [16] (see supplementary material) with the number of samples determined by the numerical bound for radius PRMs (Alg. 2) with the computed passage width and probability of failure set to $\gamma=0.1$. The connection radius required by Theorem 3.1 is ignored.
@@
We compare this strategy to two other baseline strategies. First, we compare to the sPRM motion planner with a fixed number of samples $N$ selected by hand and fixed across subproblems, which is standard practice in TAMP. Second, we compare to an incremental version of sPRM, which we call iPRM. This algorithm iteratively calls sPRM with an increasing max sample bound $K$, which increases geometrically according to $M_{i+1}=cM_{i}$ until a motion plan is found or $M\geq{N}$. Our experiments use $c=1.1$. We compare these strategies with aPRM in Fig. 4(a) for various selections of $N$.
@@
In the planar problem, we observe that aPRM is able to solve nearly all problems in less wallclock time (including numerical bound evaluation) than sPRM and iPRM. We observed that the bound calculation time was negligible with respect to the PRM planning time. The radius numerical bound helps the planner effectively trade sampling effort with the complexity of the motion planning problems, whereas sPRM either undersamples or oversamples problems with a single threshold, and iPRM wastes computation time building up to the correct number of samples. In the 6DoF problem, our results are weaker, where some instances of iPRM scale up to a sufficient number of samples faster than aPRM takes in a single computation. This is consistent with our observation that the numerical bound becomes less useful as the number of dimensions increases.
@@ 6 Related Work
Another approach to implicitly allocate samples to a sampling-based motion planner is to write an algorithm to decide the feasibility of the motion planning problem. A sampling-based motion planner can then be terminated if the motion planning problem is found to be infeasible. One common approach is to show the existence of a ‘disconnection proof,’ which consists of a hypersurface in the obstacle region that fully encloses the goal from the starting configuration [1, 19, 20]. Others use volumetric approximations of the configuration space or obstacle region [33, 25, 35] to refute feasibility (up to the accuracy of the approximation). These methods complement our approach to motion planning in TAMP. While allocating samples according to a bound (as done in Alg. 2) may help a TAMP planner solve a feasible problem with an appropriate number of samples, the planner can still increase the sample count arbitrarily in an infeasible problem after an unbounded number of attempts. Developing a suitable criterion to have the planner call an infeasibility checker before adding a large number of samples is an interesting direction for future work.
@@
Much of our understanding of PRMs with random sampling is from asymptotic results stated in the limit of infinite samples. Karaman and Frazzoli [16] proved that paths returned by PRM (and the more space-frugal PRM*) are both asymptotically optimal. Tighter estimates on the sufficient connection radius to ensure graph connectivity have been obtained as well [30]. All of these results are proven using arguments that are difficult to extend to finite samples.
@@
In addition to Hsu et al.’s work on configuration space expansiveness [12], there are several instances of finite-random sample bounds for PRMs. Kavraki et al. [17] and Dobson et al. [8] gave a bound on the number of samples required to return a feasible and near-optimal path, respectively, by assuming the clearance and path length. The stronger assumption on path length allows for simpler arguments that do not rely on covering sets, but also preclude these bounds from application in scenarios where the path length is unknown.
Deterministic sampling schemes, if found, are more sample-efficient than random samples in covering the space under many definitions of density [27]. Tsao et al. develop a fixed set that meets the dispersion density measure, the radius of the largest unsampled ball in the sample space. Likewise, for our study, we wish to have an incremental sequence that causes the dispersion to decrease as samples are added. A nonconstructive argument by Niederreiter [27] shows the existence of an optimally sample-efficient sequence for the dispersion of balls in the $\ell_{\infty}$-norm (hypercubes), which suboptimally addresses the dispersion in the $\ell_{2}$-norm related to our study. To our knowledge, an algorithm that explicitly constructs this sequence has yet to be found. Finding an optimal deterministic sequence that directly minimizes dispersion in the $\ell_{2}$-norm is related to the notoriously difficult sphere-packing problem [23, 6]. Algorithmic constructions of sequences that address other density measures exist [21, 22] but exploit regular structure available in the other density measures that is unavailable for dispersion. Empirically, these other sequences are more sample-efficient than random samples in low dimensions [15], a benefit that becomes negligible in higher dimensions [14]. For these reasons, we choose to study random sample sequences.
@@
Sample bounds for random coverings have been studied in discrete geometry and computational learning theory [11, 24, 26]. The techniques to prove these bounds have been credited to Vapnik and Chervonenkis [32], Sauer [28], and Shelah [29], who sought to find small coverings by the probabilistic method (see supplementary material). Blumer et al. build on these techniques to estimate the probability that the samples form a covering [2] to form probably approximately correct learning guarantees. Their results for the smallest $\alpha$-net (right-hand argument in Lemma 2) were confirmed to be tight up to constants by Kómlos et al. [18]. This paper connects these results to motion planning.
@@
Lastly, this paper shows the potential usefulness of a sampling bound in a search-then-sample style planner. Search-then-sample is a fairly common approach to TAMP [9, 5, 7]. These planners use ad hoc sample thresholds for motion planning, which we show can be replaced with a sample bound. In addition to setting resource limits on individual motion planning subproblems, a sample bound could guide the task-level search as a heuristic to decide which motion planning problems to attempt. Some related works along these lines include Cambon et al. [4], wherein the number of failed queries is a heuristic to decide which task plans to prioritize for motion computation.